• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name D. Parno

    Log entry time 08:01:47 on February17,2009

    Entry number 261599

    keyword=Owl shift summary 02/17

    Tuesday Owl Shift
    Shift Leader: Youcai Wang (12:00 - 2:00 a.m.)
    Diana Parno (2:00 - 8:00 a.m.)
    Shift Worker: Navaphon Muangma

    00:00 beam tuning already 30 minutes,
    continue tuning beam
    target at empty position.
    00:40 MCC asks about Magnet before target
    00:50 Ask MCC about the situation of beam tuning, they are troubleshooting the problem.
    1:10 MCC tells they need to fix the beam problem after expert comes to check.
    01:35 go to Controlled Access for experts to fix problem with Moller quads.

    07:00 Change BB T6 threshold to -20 mV and start taking cosmics data.

    Runs 1510 and 1511 are cosmics runs.

    End of shift: Beam still down, hall still in controlled access as we enter the time scheduled for beam studies.